OpenStudy (trisaba):

P(x> or = x line over + 2 standard deviation) medal and fan

OpenStudy (trisaba):

i got one more two similar

OpenStudy (trisaba):

normal distributions

OpenStudy (trisaba):

x line over means there is a line over the x

OpenStudy (trisaba):

@superman36

OpenStudy (trisaba):

@Jaynator495

OpenStudy (trisaba):

someone

OpenStudy (anonymous):

\[P(x\ge \bar{x}+2\sigma)~~?\] How is \(x\) distributed? If you're given a normal distribution, this is straightforward if you know the empirical rule: https://en.wikipedia.org/wiki/68%E2%80%9395%E2%80%9399.7_rule

OpenStudy (trisaba):

it has no labels they want like 68% is 34% + 34%

OpenStudy (trisaba):

@SithsAndGiggles

OpenStudy (trisaba):

we have to use the normal ditribution example ones from this line to this

OpenStudy (trisaba):

3 and 4

OpenStudy (trisaba):

4 and 5 sorry wrong numbers

OpenStudy (trisaba):

what am i supposed to do

OpenStudy (anonymous):

 From the rule, you have that \[\begin{cases} D+E=68\\ C+D+E+F=95\\ B+C+D+E+F+G=99.7\\ A+B+C+D+E+F+G+H=1 \end{cases}\] The distribution is symmetric, so you can determine that \[\begin{cases} D=E=34\\ C=F=13.5\\ B=G=2.35\\ A=H=0.15 \end{cases}\]